anna is seven years younger than Bradi three years from now Anna will be twice as old as brandi was nine years ago How old will Brandi be six years from now?
Answer by Maths68(1474)

Let
Present age of Anna = a
Present age of Brandi = b


Given
=====
Anna is seven years younger than Brandi
a=b-7............(1)
Three years from now
Anna will be = a+3
Nine year ago
Brandi was = b-9

Then

Anna will be twice as old as brandi was nine years ago
a+3=2(b-9)
a+3=2b-18
a=2b-18-3
a=2b-21................(2)
Put the value of a from (1) to (2)
b-7=2b-21
b-2b=-21+7
-b=-14
Multiply by -1 both sides of above equation
(-1)(-b)=(-1)(-14)
b=14
Put the value of b in (1)
a=b-7
a=14-7
a=7
Present age of Anna = a = 7 years old
Present age of Brandi = b = 14 years old
Six years from now Brandi will be = 14+6 = 20